40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Nov 19 20:09:57 2010, Zach, Laser, GYRO, realignment, etc 
    Reply  Sat Nov 20 15:02:05 2010, rana, Laser, GYRO, realignment, etc 
       Reply  Sun Nov 21 16:24:06 2010, Zach, Laser, GYRO, realignment, etc EOM_TF.pngEOM_sweep_setup.png
Message ID: 1173     Entry time: Sun Nov 21 16:24:06 2010     In reply to: 1172
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: realignment, etc 

I did the EOM sweep measurement. It looks like we are fine above ~30 MHz, but the junk at low frequencies could explain why we had excess RFAM at, say, 18 MHz. The plot and a diagram of the experimental setup are below.

EOM_TF.pngEOM_sweep_setup.png

The source level was the maximum 15 dBm, the frequency range was 1-100 MHz (linear) over 801 pts (maximum) with 30 averages. The RFAM amplitude is now back below about -80 dBm for the standard DC PD voltage of 100 mV we have been using. Note: the strange ~2 MHz oscillation for low incident power remains...

I also reconfigured the AOM double-pass setup, realigning the polarizing optics and extracting the double-first-order beam. I was able to get ~50% efficiency (~125 mW out / 250 mW in), or a single-pass efficiency of ~71%. Tomorrow I will direct this beam into the cavity and maximize transmission, then continue with the loop measurements.

 

Quote:

Remember that you still have to check the crystal for resonances. Sweep the EOM and look at the CCW PD with the cavity blocked internally. The mechanical Q of the crystal can be ~1000s, so it needs to be a fairly fine sweep.

If there are resonances close to the 33 MHz, you'll have to tune it off a little bit. And where's the sketch for the Pomona matching circuit?

 

 

ELOG V3.1.3-